# Enhancing Systematic Decompositional Natural Language Inference Using Informal Logic

### v2

@@ 4.1 RDTE Results
Our primary evaluation metric is F-score ($\beta=0.5$), meaning we put double precedence on precision over recall. Precision is particularly crucial for our needs: nontrivial false positive rates in a backward-chaining search can quickly create error propagation, wasted time on invalid search branches, and worse trees. Contrastly, while recall is important, it is less catastrophic to overfilter valid decompositions than to underfilter bad ones.
[…]
We also find that knowledge distillation proves highly effective on RDTE: the student ChatjGPT models improve 5-13% over the closest ChatGPT methods, while the fine-tuned RoBERTa student ultimately outperforms the teacher GPT-4 method on both datasets. Table 4 shows that this is the result of substantially higher precision (68 to 55) at the expense of recall (57 to 90).
Figure 3: TreeWise generates many premise decompositions of a hypothesis and checks whether any candidates are valid entailments. Premises are then recursively decomposed until it finds any tree(s) fully grounded, via entailment, in one or more documents from a corpus like Wikipedia. Statements entailed by a document are generated via forward chaining, while the rest of the search is backward. TreeWise overgenerates enough decompositions that many end up untraversed due to the search budget or nonentailment.
@@ 5 TreeWise
To illustrate the impact that an RDTE-based entailment model can have on systematic reasoning, we apply it as a module in a new, state-of-the-art entailment engine called TreeWise: Textual Reasoning Engine with Enriched Ways to Intelligently Search for Entailment. TreeWise builds upon the backward chaining entailment tree search framework first introduced by Nellie Weir et al. (2023). The core functionality of TreeWise is to answer the question, “is NL hypothesis $H$ compositionally entailed by a corpus of documents $C$ in the context of a question $Q$?” Illustrated in Figure 3, the search algorithm attempts to ground the hypothesis via decomposition into premises entailed by passages in a verified corpus such as Wikipedia, performing a recursive search over candidate trees. Implemented in Prolog, the algorithm follows a breadth-first strategy, performing the following at each recursive hypothesis $H$:
@@
    Decompose $H$ into a set of $d$ candidate decompositions using a series of prompts to the LLM. We use a heterogeneous set of prompts to propose a diverse set of candidates that encourage using the $i$ inferences.
@@
    Condense the decompositions by identifying paraphrases and deduplicating semantically equivalent candidates.
@@
This process is repeated until either a user-specified expansion budget is exhausted, or a proof is found and no search branches remain that could yield a higher score. The system is designed to be amenable to a costly API; it batches together decompositions of a single hypothesis into one prompt, then asks it to score them all at once. This is a crucial optimization, as it allows us to score hundreds of decompositions in a single API call.
TreeWise represents a substantial overhaul of the T5-based Nellie system, including incorporating a variety of prompting methods, forward chaining, and the capacity to reason over longer passages. We include an overview of search logic and full architectural comparison in Appendix A.
@@ 6.1 QA Evaluation
We run a series of QA experiments to highlight the effectiveness of applying RDTE-trained distillation student models to entailment tree creation algorithms. This framework suggests a way forward for improving entailment engines in new domains: first extract non-optimized reasoning traces from the engine, annotate them under the RDTE protocol with GPT-4 (which is itself too slow and expensive to use in the engine), train student models on the silver data, and then substitute the student as an entailment classifier in the engine. The experiments below emulate this scenario and show that the resulting engines improve on both QA and on generated tree quality.
Datasets and Scenarios
We evaluate TreeWise and a series of entailment tree-creating baselines on the two QA datasets matching those used to construct RDTE: 340 ARC questions from the EntailmentBank test set (EBQA), and 419 HotpotQA questions recast as multiple choice by using GPT-4 to generate incorrect answer options.
@@
Metrics We take a two-pronged approach to evaluating QA systems: they should produce both strong end-task accuracy while also producing coherent and logically sound entailment trees explaining the chosen answer. To evaluate the latter, we introduce a new model-based tree integrity score. To score an entailment tree, we slice it up into its component entailment steps and then use GPT-4 to score them under the RDTE protocol. Following the intuition that an argument is only as strong as its weakest link, we take the minimum such score as the tree’s overall integrity score.
Baselines We compare a version of TreeWise using the RDTE-trained student ChatGPT and RoBERTa models to an identical version that uses an ICL prompted ChatGPT and the RoBERTa model used by TreeWise, which is trained on non-RDTE entailment data. We also use ChatGPT as the decomposition generator and QA2D model for TreeWise, meaning it does not leverage GPT-4 for anything except knowledge distillation. We also evaluate a set of greedy baselines for entailment tree creation. These baselines are designed to mimic the behavior of TreeWise in a simplified manner without the systematic search algorithm:
@@
Psuedocode for these is provided in §2. Each process is repeated 5 times, yielding 5 different candidate trees for each answer choice, then fed to the tree integrity scorer using the student ChatGPT to score each tree. We take the highest-scoring tree and corresponding answer as the final output. We compare these to versions where the student ChatGPT is replaced by regular ChatGPT.
Results
Table 3 shows QA results for these methods. We observe that for all methods, tree integrity score increases when using the knowledge distilled student. In all cases but 1 baseline on HotpotQA, QA accuracy also goes up by 1 to 7 points. We observe that TreeWise achieves the highest QA and integrity scores, while the stepwise outperforms the end-to-end generator under integrity but vice versa for two QA scenarios. See Figure 13 and 14 for example trees generated by TreeWise using Wikipedia as its knowledge source.
@@ 7 Related Work
Neuro-Symbolic Search Algorithms Numerous recent papers have explored algorithms for constructing entailment trees given a set of support facts Dalvi et al. (2021); Ribeiro et al. (2023b); Yang et al. (2022); Bostrom et al. (2022). There is also growing literature on backward- and forward-chaining algorithms for performing entailment tasks Creswell et al. (2023); Weir et al. (2023). Previous generations of systems have also explored entailment search algorithms, e.g. NaturalLI Angeli and Manning (2014).
Annotating Textual Entailment Datasets The PASCAL RTE Challenge Dagan et al. (2005), SNLI Bowman et al. (2015) and MNLI Williams et al. (2018) have led to numerous NLI leaderboards. These annotation efforts (A) are generally not tied to downstream complex reasoning tasks; and (B) commonly target uncontroversial items for which annotators have high agreement without requiring granular task specifications. The recent BaRDa dataset Clark et al. (2023) throws out a large fraction of collected decomposition annotations because of low agreement. In contrast, when constructing RDTE, we specifically target the hard-to-evaluate items that unavoidably manifest during systematic reasoning algorithms. We find that state-of-the-art LLMs like GPT-4 achieve less than 60% F-score on the new challenge set. Our use of ordinal annotation draws from JOCI Zhang et al. (2017), a collection of common sense inferences annotated on a 5-point scale of likeliness.
Annotator Disagreement Various works have come up against the challenge of annotator disagreement for reasoning tasks. For example, AmbiFC Glockner et al. (2021), concerned with whether a piece of evidence supports a given claim, explores the common factors causing annotator disagreements. Pavlick and Kwiatkowski (2019) found persistent patterns of annotator disagreement amongst RTE judgments not simply due to noise; efforts such as the UNLI protocol Chen et al. (2020) aim to address this concern by modeling annotations on a logistic probability scale.
Computational Argumentation Existing work has explored NLP methods for computational argumentation; one such subfield is argumentation mining Palau and Moens (2009), which aims to detect and relate the arguments in a text passage. Jin et al. (2022) collect a dataset of items exhibiting 14 different fallacies. Stab and Gurevych (2017), in similar spirit to our work, hand-annotate the sufficiency of argumentative essay sections using Blair and Johnson’s SAR criteria, but is not geared towards the decompositional entailment we consider.
Assessing Explanations Our work builds on a growing body of literature on evaluating the role and quality of model-generated explanations Wiegreffe and Marasovic (2021); Tan (2022). Our paper fits into this broader discourse by specifically investigating how to improve the quality and effectiveness of decompositional entailment items that drive reasoning in the context of complex tasks. Similar to our work is Valentino et al. (2021), who introduce the EEV methodology by which an explanation is translated into logical form and then assessed for quality by a trained semanticist. Our RDTE approach diverges by focusing on the decompositional textual entailment specifically in the context of tree-based reasoners for which the generated explanation undergirds an explicit decision process (i.e. a search). While EEV depends on translation into formal logic, we construct an informal logic-based framework that does not require a trained semanticist to annotate; we also demonstrate the shortcomings of existing models at assessing validity, and also demonstrate practical applications in improving reasoning engines like TreeWise through knowledge distillation.
@@ 8 Conclusion
The trustworthy application of LLMs to complex reasoning tasks critically depends not only on accurate responses, but also requires accurate justifications. To date, research in proof-backed reasoning has largely focused on measuring response accuracy, on the faulty presumption that an accurate response implies a correspondingly accurate proof.
In this work we develop a protocol for the assessment of compositional entailment, based on rubric derived from works in informal logic. This protocol was employed in an annotation procedure resulting in a novel, high quality dataset of judged compositional entailments called RDTE, along with tens of thousands of automatically scored items by GPT-4 under this rubric, in multiple domains. Further, we demonstrated this work supports state-of-the-art results, when coupled to a novel system for evidence-grounded entailment-tree generation. Of note, we are the first to include a non-scientific reasoning domain as a target for experimentation, demonstrating our success on HotPot.
@@ A.1 Search Logic
    $\textsc{Prove}(h)\Leftarrow\textsc{RuleGen}(h^{+},f_{1}^{-},f_{2}^{-})\ \land$ $\quad\textsc{Entails}([f_{1},f_{2}],h)$ $\land\ \textsc{Prove}(f_{1})\land\textsc{Prove}(f_{2})$
@@
    $\textsc{Prove}(h)\Leftarrow\textsc{Retrieve}(h^{+},f_{1}^{-})\ \land$ $\quad\textsc{RuleGen}(h^{+},f_{1}^{+},f_{2}^{-})$ $\ \land\ \textsc{Entails}([f_{1},f_{2}],h)\ \land$ $\quad\textsc{Prove}(f_{2})$
@@
To address these issues, we make the following modifications, replacing rules 2 and 3 with rules 4, $*{}^{*}$, $*{}^{*}$. The latter two rules are only executed if 4 does not succeed. Bolded symbols denote string lists.
@@
    $\textsc{Prove}(h)\Leftarrow\textsc{RuleGen}(h^{+},[],\mathbf{f})\ \land$ $\quad\textsc{Entails}(\mathbf{f},h)$ $\land\ \textsc{MapList}(\mathbf{f}^{+},\textsc{Prove})$  [^9] https://www.swi-prolog.org/pldoc/man?predicate=maplist/2
  • ${}^{\ast}$.
@@
    $\textsc{Prove}(h)\Leftarrow\textsc{Retrieve}(h^{+},\mathbf{s}^{-})\ \land$ $\quad\textsc{InferenceGen}(h^{+},\mathbf{s}^{+},\mathbf{i}^{-})\ \land$ $\quad\textsc{RuleGen}(h^{+},\mathbf{i}^{+},\mathbf{f}^{-})\ \land\ \textsc{Entails}(\mathbf{f},h)\ \land$
@@ A.4 Reasoning Filters
Our RDTE-oriented prompting strategy discussed in the main body of the paper is shown in Figure 6, 7, and 8. The zero-shot variant contains the same directions but no exemplars. We use a separate set of exemplars for Hotpot vs ARC.
@@ Appendix C Example TreeWise Outputs
Factuality
How factual is the fact? 1 is completely false, 5 is completely true.
Relevancy
How relevant is the fact to helping decompose the conclusion? An irrelevant fact is one that either does not address a key aspect of the conclusion, introduces irrelevant information, or is otherwise unnecessary should be scored lower in relevance. 1 is completely irrelevant, 5 is completely relevant. If the fact is situationally relevant to the conclusion, but contradicts it, you should still score it 5.
Redundancy
Does the fact introduce new information that is not already contained in other facts in the decomposition? 1 is completely redundant, e.g., the third fact completely restates the first one, and 5 is completely new information. Sometimes, one of the facts directly restates the entire conclusion by itself, which should be marked with the checkmark only and should not affect the numerical score. Otherwise, facts including information included in the conclusion are acceptable and strictly necessary.
@@
Is a fact ambiguously grounded in the question context in a way that does not affect the reasoning? (e.g. the fact “two sticks are rubbed together” in the question “what is an example of a force producing heat? (A) two sticks rubbed together, …”)
This is acceptable and can be 5/5 factuality
Is a fact true in nearly all cases except extreme ones that don’t pertain to the question?
(Yes = 5 factuality)
@@
Is a fact not on topic? (“on topic” is defined as containing nouns or entities that appear in the hypothesis)
(Yes = 1 relevancy)
Does there not exist some (potentially over-pedantic) decomposition in which the given fact would be necessary to complete the entailment?
(Yes = max 2 relevancy)
Would removing an on-topic fact in isolation not change the extent to which the conclusion is supported?
(Yes = 2 relevancy)
Would removing an on-topic fact in isolation minimally change the extent to which the conclusion is supported?
(Yes = 3 relevancy)
@@
Is a fact a paraphrase of another fact?
(Yes = 1 redundancy
@@
Does a given fact add entailment in isolation, but if you removed the fact conditioned on the rest of the facts, it would not change the extent to which the conclusion is supported?
(Yes = max 2 redundancy)
Does a fact restate information in the question text (not the conclusion)?
This is acceptable. Only check for restatements of other facts and/or the conclusion. Restatement of the question text, especially to cite evidence, is fine.
@@
2 (Poor)
Any two of the following: (1) some nontrivial amount of redundancy, (2) one irrelevant fact (2/5 or lower), (3) missing/implicit information that makes deducing the conclusion impossible without a substantial leap in logic. Would not convince a human of the conclusion.
@@
5 (Perfect)
Completely correct and sound decomposition. No redundancy and no missing/implicit information. No ambiguous language. Addresses all conditions required to infer the conclusion. Does not leave anything implicit.
@@
Are any of the premises not well-formed? (no fragments, no 1 sentence that was split into two non-sentence parts)
(Yes = 1)
Do the premises together or individually contradict the conclusion instead of supporting it?
(Yes = 1)
@@
Are there between-premise pronoun references (‘this’, ‘that’, ‘such’) whose antecedent would be ambiguous without the other premises?
(Yes = 1)
Are there any conjunctive adverbs like “therefore" or “thus"?
(Yes = 1)
@@
Are all premises irrelevant, off-topic, or not contributing any correct logic? (e.g. all 1’s for relevance, or removing all of them would not change the extent of the entailment)
(Yes = 1)
@@
Does any premise essentially restate the conclusion without adding/removing any information?
(Yes = max 2)
@@
Are there at least two of the following? (1) redundant fact, (2) untrue fact, (3) irrelevant fact, (4) missing information
(Yes = max 2)
@@
Is the conclusion assuming an effect that isn’t directly linked to the cause, or overlooking more immediate effects?
(Yes = max 2)
@@
If you removed all redundant or irrelevant (3/5 or lower) facts, would there be only one fact remaining and not full entailment?
(Yes = max 2)
Is there any amount of logical error present?
(Yes = max 2)
Did you give any fact a 2/5 or lower for factuality or relevance?
(Yes = max 3)
Does proving one premise amount to proving all of the others?
(Yes = max 3)
@@
Are the premises all factual and relevant, but there is a part of the conclusion (e.g. something non-common-sense or a thing that a 10-year-old would not intuit in the context of the question) that is not stated or explained?
(Yes = max 3)
@@
If you removed all redundant or irrelevant (3/5 or lower) facts, would there be only one fact remaining?
(Yes = max 3)
@@
Are the premises all evidence statements entailed by the question context and nothing else?
(Yes = likely max 3)
Is one of the facts not true, but if it were then it’d be a perfect entailment?
(Yes = 3)
Did you give any fact a 3/5 or lower for redundancy, factuality, or relevance?
(Yes = max 4)
@@
Are there two separate arguments being partially/mostly made to support the hypothesis, but one/both is missing some implicit premises?
(Yes = max 4)
@@
Are the premises all factual and relevant, but the conclusion does not follow from the premises for a minor reason (e.g. a common sense-y fact that would have been inferred by a 10-year-old in the context of the question)
(Yes = max 4)
Do two of the facts perfectly entail the conclusion, but the third is essentially redundant?
(Yes = 4)
Is the conclusion irrelevant to the question, but the entailment supports the conclusion perfectly?
(Yes = 5)
@@
Does the question ask something along the lines of “which is the best…?” and the entailment doesn’t mention the other answer options?
Treat it like the “best” is not there
@@
Do the premises not properly entail the conclusion for some other reason?
Reach out to us for clarification
@@
Is one premise P1 an effective paraphrase of the hypothesis H, but another premise P2 serves as lexical grounding between P1 and H?
Ignore the paraphrase = max 2 rule
@@
Document-conditioned forward generation prompt:
@@
{{"index": 5, "factuality": [5, 2, 5], "relevance": [3, 3, 5], "redundancy": [1, 5, 5], "complete_inference": 3, "explanation": "Not all objects conduct electricity, so the premise ’objects conduct electricity’ is an overgeneralization."}}
@@
Document-conditioned forward generation prompt (continued):
@@
(2) Competition is a natural occurrence when resources are limited AND Unlimited resources can lead to an absence of competition AND Without competition, there’s no natural selection

### v3

@@ 4.1 RDTE Results
Our metric is F-score ($\beta=0.5$), putting double precedence on precision over recall. Precision is particularly crucial for our needs: false positives in a backward-chaining search can quickly create error propagation, wasted time on invalid search branches, and worse trees. Contrastly, while recall is important, it is less catastrophic to overfilter valid decompositions than to underfilter bad ones.
Table 1 shows F0.5 performance by models on the RDTE dataset. We also display precision and recall in Table 3. We observe that no model cracks 70, suggesting room for future improvement. Overall we find the RDTE-oriented prompts to GPT-4 outperform all existing methods for compositional entailment by 10 or more points. [^9] Table 1 shows GPT-4’s score when thresholding on 5, not 4, which we found to score lower (see Table 3).
We find that knowledge distillation proves effective on RDTE: the student ChatGPT models improve 5-13% over the closest ChatGPT methods, while the fine-tuned RoBERTa student ultimately outperforms the teacher GPT-4 method. Table 3 shows that this is due to higher precision (68 to 55) at the expense of recall (57 to 90). The RoBERTa student is the most precise classifier tested, and the ChatGPT student is the next-highest among non-GPT-4 models; however, both show much lower recall than the others tested.
Figure 4: TreeWise generates many premise decompositions of a hypothesis and checks whether any candidates are valid entailments. Premises are then recursively decomposed until it finds any tree(s) fully grounded in one or more documents from a corpus like Wikipedia. Statements entailed by documents are generated via forward chaining, while the rest of the search is backward. Many decompositions end up untraversed due to the search budget or nonentailment.
@@ 5 TreeWise
To illustrate the impact that an RDTE-based entailment model can have on systematic reasoning, we apply it as a module in a new, state-of-the-art entailment engine called TreeWise: Textual Reasoning Engine with Enriched Ways to Intelligently Search for Entailment. TreeWise builds upon the backward chaining entailment tree search framework first introduced by Nellie Weir et al. (2023). The core functionality of TreeWise is to answer the question, “is NL hypothesis $H$ compositionally entailed by a corpus of documents $C$ in the context of a question $Q$?” Illustrated in Figure 4, the search attempts to ground the hypothesis via recursive decomposition into premises entailed by passages in a verified corpus such as Wikipedia. Implemented in Prolog, the algorithm follows a breadth-first strategy, performing the following at each recursive hypothesis $H$:
@@
    Decompose $H$ into a set of $d$ candidate decompositions using a series of prompts to the LLM. We use a diverse set of prompts to propose candidates that encourage using the $i$ inferences.
@@
    Condense the decompositions by deduplicating semantically equivalent candidates.
@@
This process is repeated until either a user-specified expansion budget is exhausted, or a proof is found and no search branches remain that could yield a higher score. The system is designed to be amenable to a costly API; it batches together decompositions of a single hypothesis into one prompt, then asks it to score them all at once.
TreeWise represents an overhaul of the T5-based Nellie system, incorporating a variety of prompting methods, forward chaining, and the capacity to reason over long passages. We discuss the search logic and architectural comparison in §A.
@@ 6 TreeWise Experiments
The distillation framework suggests a way forward for improving entailment engines in new domains: first extract non-optimized reasoning traces from the engine, annotate them under the RDTE protocol with GPT-4 (which is itself too slow and expensive to use in the engine), train student models on the silver data, and then substitute the student as an entailment classifier in the engine. The experiments below emulate this scenario and show that the resulting engines improve on both QA and on generated tree quality, highlighting the effectiveness of applying RDTE-trained distillation student models to entailment tree creation algorithms.
Entailment Tree QA Task Definition We pursue a version of evidence-grounded explainable QA in which models are not only right, but right for the right reasons. Given corpus $C$, multiple-choice question $Q$ and hypotheses $h_{1},\dots,h_{m}$ corresponding to answer options $a_{1},\dots,a_{m}$, we task models to predict correct answer $a_{i}^{*}$ and generate entailment tree $T_{h_{i}}$ rooted at $h_{i}$ such that (1) the $T_{h_{i}}$’s leaves are in $C\cup Q$ and (2) the tree’s steps explain how each parent follows compositionally from its children. Models are evaluated on selecting $a_{i}^{*}$ as well as the correctness of the reasoning in $T_{h_{i}}$.
@@ 6.1 QA Evaluation
Datasets We evaluate TreeWise and a series of entailment tree-creating baselines on the two QA datasets used to construct RDTE: 340 test questions from the EntailmentBank (EBQA), and 419 HotpotQA questions recast as multiple choice by using GPT-4 to generate incorrect answer options.
@@
Metrics We take a two-pronged approach to evaluating systems: they should produce both strong QA accuracy and coherent and logically sound entailment trees explaining the chosen answer. To evaluate the latter, we introduce a new model-based tree integrity score. We use GPT-4 to score each of a tree’s component entailment steps under the RDTE protocol. Following the intuition that an argument is only as strong as its weakest link, we take the minimum score as the tree’s overall score.
Baselines We compare a version of TreeWise using the RDTE-trained student ChatGPT and RoBERTa models to an identical version that uses an ICL prompted ChatGPT and the RoBERTa model used by Nellie, which is trained on non-RDTE entailment data. We use ChatGPT as the decomposition generator and QA2D model for TreeWise, meaning it does not use GPT-4 for anything at test time. We also evaluate a pair of greedy baselines for entailment tree creation. These baselines are designed to mimic TreeWise’s behavior in a simplified manner without the systematic search:
@@
Psuedocode for these is provided in §2. Each process is repeated 5 times, yielding 5 different candidate trees for each answer choice, then fed to the tree integrity scorer using the student ChatGPT to score each tree. We take the highest-scoring tree and corresponding answer as the final output. We compare these to non-RDTE versions where the student ChatGPT is replaced by regular ChatGPT.
Results Table 2 shows QA results for these methods. We observe that for all methods, tree integrity score increases when using the knowledge distilled student. For all but 1 baseline on HotpotQA, QA accuracy also goes up by 1 to 7 points. We observe that TreeWise achieves the highest QA and integrity scores, while the stepwise outperforms the end-to-end generator under integrity but vice versa for two QA scenarios. See Figure 15 and 16 for example trees generated by TreeWise using Wikipedia as its knowledge source.
@@ 7 Related Work
Annotating Textual Entailment Datasets The PASCAL RTE Challenge Dagan et al. (2005), SNLI Bowman et al. (2015) and MNLI Williams et al. (2018) have led to numerous NLI leaderboards. These datasets (A) are generally not tied to downstream reasoning tasks; and (B) commonly target uncontroversial items for which annotators have high agreement without needing granular instructions. The BaRDa dataset Clark et al. (2023) throws out a large fraction of decomposition annotations because of low agreement. In contrast, RDTE specifically targets the hard-to-evaluate items that unavoidably manifest during systematic reasoning algorithms. Our use of ordinal annotation draws from JOCI Zhang et al. (2017), a collection of common sense inferences annotated on a 5-point scale of likeliness.
Annotator Disagreement Various works have faced the challenge of annotator disagreement for reasoning tasks. AmbiFC Glockner et al. (2021), concerned with whether a piece of evidence supports a given claim, explores the common factors causing disagreements. Pavlick and Kwiatkowski (2019) found persistent patterns of disagreement amongst RTE judgments not simply due to noise; efforts such as the UNLI protocol Chen et al. (2020) aim to address this concern by modeling annotations on a logistic probability scale.
Computational Argumentation Existing work has explored methods for computational argumentation; one subfield is argumentation mining Palau and Moens (2009), which aims to detect and anaylze the arguments in a passage. Jin et al. (2022) collect a dataset of items exhibiting 14 types of fallacies. Stab and Gurevych (2017) hand-annotate the sufficiency of argumentative essays using RAS criteria, but are not geared towards entailment.
Assessing Explanations Our work builds on literature evaluating the role and quality of model-generated explanations Wiegreffe and Marasovic (2021); Tan (2022). We add to this discourse by investigating how to improve the quality of decompositional entailment items that drive reasoning for complex tasks. Similar to ours is the EEV methodology Valentino et al. (2021), under which an explanation is translated into logical form then assessed by a trained semanticist. RDTE requires neither the translation nor the trained semanticist.
@@ 8 Conclusion
The trustworthy application of LLMs to complex reasoning tasks critically depends not only on accurate responses, but also accurate justifications. To date, research in entailment tree reasoning has largely focused on measuring the former.
This work develops a protocol based on works in informal logic for the assessment of compositional entailment, the core reasoning task undergirding entailment trees. We employed this rubric-based protocol to annotate RDTE, a high quality dataset of judged compositional entailments, along with tens of thousands of automatically scored items by GPT-4 under the same instructions, in multiple domains. We demonstrated that RDTE supports state-of-the-art results by a novel system for evidence-grounded entailment tree generation.
@@ A.1 Search Logic
    $\textsc{Prove}(h)\Leftarrow\textsc{RuleGen}(h^{+},f_{1}^{-},f_{2}^{-})\ \land$
    $\quad\textsc{Entails}([f_{1},f_{2}],h)$ $\land\ \textsc{Prove}(f_{1})\land\textsc{Prove}(f_{2})$
@@
    $\textsc{Prove}(h)\Leftarrow\textsc{Retrieve}(h^{+},f_{1}^{-})\ \land$
    $\quad\textsc{RuleGen}(h^{+},f_{1}^{+},f_{2}^{-})$ $\ \land\ \textsc{Entails}([f_{1},f_{2}],h)\ \land$
    $\quad\textsc{Prove}(f_{2})$
@@
To address these issues, we make the following modifications, replacing rules 2 and 3 with rules 4, 2, 3. The latter two rules are only executed if 4 does not succeed. Bolded symbols denote string lists.
@@
    $\textsc{Prove}(h)\Leftarrow\textsc{RuleGen}(h^{+},[],\mathbf{f})\ \land$
    $\quad\textsc{Entails}(\mathbf{f},h)$ $\land\ \textsc{MapList}(\mathbf{f}^{+},\textsc{Prove})$  [^11] https://www.swi-prolog.org/pldoc/man?predicate=maplist/2
  • 3.
@@
    $\textsc{Prove}(h)\Leftarrow\textsc{Retrieve}(h^{+},\mathbf{s}^{-})\ \land$
    $\quad\textsc{InferenceGen}(h^{+},\mathbf{s}^{+},\mathbf{i}^{-})\ \land$
    $\quad\textsc{RuleGen}(h^{+},\mathbf{i}^{+},\mathbf{f}^{-})\ \land\ \textsc{Entails}(\mathbf{f},h)\ \land$
@@ A.4 Reasoning Filters
Our RDTE-oriented prompting strategy discussed in the main body of the paper is shown in Figure 7, 8, and 9. The zero-shot variant contains the same directions but no exemplars. We use a separate set of exemplars for Hotpot vs ARC.
@@ Appendix C Example TreeWise Outputs
Factuality How factual is the fact? 1 is completely false, 5 is completely true.
Relevancy How relevant is the fact to helping decompose the conclusion? An irrelevant fact is one that either does not address a key aspect of the conclusion, introduces irrelevant information, or is otherwise unnecessary should be scored lower in relevance. 1 is completely irrelevant, 5 is completely relevant. If the fact is situationally relevant to the conclusion, but contradicts it, you should still score it 5.
Redundancy Does the fact introduce new information that is not already contained in other facts in the decomposition? 1 is completely redundant, e.g., the third fact completely restates the first one, and 5 is completely new information. Sometimes, one of the facts directly restates the entire conclusion by itself, which should be marked with the checkmark only and should not affect the numerical score. Otherwise, facts including information included in the conclusion are acceptable and strictly necessary.
@@
Is a fact ambiguously grounded in the question context in a way that does not affect the reasoning? (e.g. the fact “two sticks are rubbed together” in the question “what is an example of a force producing heat? (A) two sticks rubbed together, …”) This is acceptable and can be 5/5 factuality
Is a fact true in nearly all cases except extreme ones that don’t pertain to the question? (Yes = 5 factuality)
@@
Is a fact not on topic? (“on topic” is defined as containing nouns or entities that appear in the hypothesis) (Yes = 1 relevancy)
Does there not exist some (potentially over-pedantic) decomposition in which the given fact would be necessary to complete the entailment? (Yes = max 2 relevancy)
Would removing an on-topic fact in isolation not change the extent to which the conclusion is supported? (Yes = 2 relevancy)
Would removing an on-topic fact in isolation minimally change the extent to which the conclusion is supported? (Yes = 3 relevancy)
@@
Is a fact a paraphrase of another fact? (Yes = 1 redundancy
@@
Does a given fact add entailment in isolation, but if you removed the fact conditioned on the rest of the facts, it would not change the extent to which the conclusion is supported? (Yes = max 2 redundancy)
Does a fact restate information in the question text (not the conclusion)? This is acceptable. Only check for restatements of other facts and/or the conclusion. Restatement of the question text, especially to cite evidence, is fine.
@@
2 (Poor) Any two of the following: (1) some nontrivial amount of redundancy, (2) one irrelevant fact (2/5 or lower), (3) missing/implicit information that makes deducing the conclusion impossible without a substantial leap in logic. Would not convince a human of the conclusion.
@@
5 (Perfect) Completely correct and sound decomposition. No redundancy and no missing/implicit information. No ambiguous language. Addresses all conditions required to infer the conclusion. Does not leave anything implicit.
@@
Are any of the premises not well-formed? (no fragments, no 1 sentence that was split into two non-sentence parts) (Yes = 1)
Do the premises together or individually contradict the conclusion instead of supporting it? (Yes = 1)
@@
Are there between-premise pronoun references (‘this’, ‘that’, ‘such’) whose antecedent would be ambiguous without the other premises? (Yes = 1)
Are there any conjunctive adverbs like “therefore" or “thus"? (Yes = 1)
@@
Are all premises irrelevant, off-topic, or not contributing any correct logic? (e.g. all 1’s for relevance, or removing all of them would not change the extent of the entailment) (Yes = 1)
@@
Does any premise essentially restate the conclusion without adding/removing any information? (Yes = max 2)
@@
Are there at least two of the following? (1) redundant fact, (2) untrue fact, (3) irrelevant fact, (4) missing information (Yes = max 2)
@@
Is the conclusion assuming an effect that isn’t directly linked to the cause, or overlooking more immediate effects? (Yes = max 2)
@@
If you removed all redundant or irrelevant (3/5 or lower) facts, would there be only one fact remaining and not full entailment? (Yes = max 2)
Is there any amount of logical error present? (Yes = max 2)
Did you give any fact a 2/5 or lower for factuality or relevance? (Yes = max 3)
Does proving one premise amount to proving all of the others? (Yes = max 3)
@@
Are the premises all factual and relevant, but there is a part of the conclusion (e.g. something non-common-sense or a thing that a 10-year-old would not intuit in the context of the question) that is not stated or explained? (Yes = max 3)
@@
If you removed all redundant or irrelevant (3/5 or lower) facts, would there be only one fact remaining? (Yes = max 3)
@@
Are the premises all evidence statements entailed by the question context and nothing else? (Yes = likely max 3)
Is one of the facts not true, but if it were then it’d be a perfect entailment? (Yes = 3)
Did you give any fact a 3/5 or lower for redundancy, factuality, or relevance? (Yes = max 4)
@@
Are there two separate arguments being partially/mostly made to support the hypothesis, but one/both is missing some implicit premises? (Yes = max 4)
@@
Are the premises all factual and relevant, but the conclusion does not follow from the premises for a minor reason (e.g. a common sense-y fact that would have been inferred by a 10-year-old in the context of the question) (Yes = max 4)
Do two of the facts perfectly entail the conclusion, but the third is essentially redundant? (Yes = 4)
Is the conclusion irrelevant to the question, but the entailment supports the conclusion perfectly? (Yes = 5)
@@
Does the question ask something along the lines of “which is the best…?” and the entailment doesn’t mention the other answer options? Treat it like the “best” is not there
@@
Do the premises not properly entail the conclusion for some other reason? Reach out to us for clarification
@@
Is one premise P1 an effective paraphrase of the hypothesis H, but another premise P2 serves as lexical grounding between P1 and H? Ignore the paraphrase = max 2 rule
@@
RDTE judgment prompt:
@@
{{"index": 5, "factuality": [5, 2, 5], "relevance": [3, 3, 5], "redundancy": [1, 5, 5], "complete_inference": 3, "explanation": "Not all objects conduct electricity, so the premise objects conduct electricity is an overgeneralization."}}
@@
RDTE judgment prompt (continued):
@@
(2) Competition is a natural occurrence when resources are limited AND Unlimited resources can lead to an absence of competition AND Without competition, theres no natural selection